katrina can design a 3 feet by 3 feet board using egg shells in 3 hours.her cousin,pia volunteered to help katrina.if they work together,it will take them 1 hour and 40 minutes to finish the job.how long does it take pia to do the job done?
Answer by nerdybill(7384)

katrina can design a 3 feet by 3 feet board using egg shells in 3 hours.her cousin,pia volunteered to help katrina.if they work together,it will take them 1 hour and 40 minutes to finish the job.how long does it take pia to do the job done?
.
Let x = time (minutes) it takes pia alone
then
time it takes for both:
1 hour and 40 minutes = 100 minutes
.
katrina's rate: 1/180
.
100(1/180 + 1/x) = 1
multiplying both sides by 180x:
100(x + 180) = 180x
100x + 18000 = 180x
18000 = 80x
225 minutes = x
or
3 hours and 45 minutes = x (pia doing it alone)